# Extreme events and power-law distributions from nonlinear quantum dissipation

### v1

@@ Abstract
Power-law probability distributions are widely used to model extreme statistical events in complex systems, with applications to a vast array of natural phenomena ranging from earthquakes to stock market crashes to pandemics. We propose the emergence of power-law distributions as a generic feature of quantum systems with strong nonlinear dissipation. We introduce a prototypical family of quantum dynamical systems with nonlinear dissipation, and prove analytically the emergence of power-law tails in the steady state probability distribution for energy. The power law physically originates from the amplification of quantum noise, where the scale of the microscopic fluctuations grows with the energy of the system. Our model predicts a power-law regime with infinite mean energy, which manifests as extreme events and divergences in the measurement statistics. Furthermore, we provide numerical evidence of power-law distributions for a general class of nonlinear dynamics known as quantum Liénard systems. This phenomenon can be potentially harnessed to develop extreme photon sources for novel applications in light-matter interaction and sensing.
@@
Introduction.— The natural occurrence of power-law probability distributions underpin many fascinating statistical phenomena in complex systems across vast domains such as physics, biology, geology, economics and social science [1, 2, 3, 4]. Power-law distributions are often associated with the presence of extreme or ‘black swan’ events, which are statistical events that occur with low frequency but yet have enormous impact [5, 6]. This is sometimes stated in various contexts as the Pareto principle: a majority of the observed effects arise from a minority of the causes. Extreme events are abound in nature, with well-known examples including earthquakes [7], rogue ocean waves [8], solar flares [9], stock market crashes [10] and pandemics [11].
@@
Often, extreme events are observed as emergent phenomena in large-scale systems, and statistical models are constructed to explain the empirical data a posteriori. Here, we show that such a phenomenon can emerge dynamically in one of the most fundamental stochastic system in nature — a single dissipative quantum system. To this end, we construct a prototypical model of a quantum dynamical system with nonlinear dissipation, and prove analytically that the steady state probability distribution for the energy exhibits a power-law tail. If we regard the system physically as a quantum particle trapped in a potential well, this phenomenon manifests as a power-law decay in the position and momentum distribution of the quantum particle [12, 13].
@@
The physical origin of the power-law distribution in our model can be attributed to quantum noise [14]. The phenomenon stems from the intimate connection between microscopic fluctuations and dissipation in quantum mechanics, and traces back to Heisenberg’s uncertainty principle. From the perspective of open quantum systems, the microscopic interactions between the system and the external bath, which is responsible for the dissipation, must inject noise into the system [14, 15, 16]. Furthermore, the form of this quantum noise is highly dependent on the type of dissipation. For nonlinear dissipation, this noise is typically multiplicative, which means the noise is coupled to the quantum state of the system [17, 18]. In our model, the coupling induces a self-amplification of the noise, in which the scale of the fluctuations grow with the excitation of the system. The amplified fluctuations at large excitation are only weakly suppressed by the dissipation, and the net result is a power-law distribution.
@@
Our model predicts a regime where the power-law tail vanishes so slowly such that the quantum state, while normalizable, contains an infinite number of excitation quanta on average. Moreover, the distribution for the excitation number is highly skewed towards low excitation. This gives rise to extreme events in which the system, when measured in the excitation number basis, produces highly excited states containing orders of magnitude more excitation quanta than the typically measured values. This regime of the power law was previously observed experimentally in a different nonlinear quantum optical setup, which works by amplifying the vacuum fluctuations [19]. In that experiment, extreme emission events containing $\sim 10^{6}$ photons in a single pulse were observed.
@@
Applying our model to the context of quantum optics, where the excitations are now photons, our analytical results show that the system can exhibit an extreme case of photon superbunching whereby the emitted photons are significantly more correlated in time as compared to thermal light sources [20]. The intensity correlation function (at zero delay) in this case can be made arbitrarily high, due to the strong fluctuations in the electromagnetic field. We establish quantitative predictions for various physical quantities, which can serve as experimental signatures of the power-law distribution.
Finally, we propose that the emergence of a power-law tail is a generic feature of quantum systems with strong nonlinear dissipation. To support this claim, we study a wide range of nonlinear dissipative models proposed in Refs. [21, 22], which are quantum analogs of the Liénard system. The Liénard system is a general class of nonlinear dynamical models containing many paradigmatic examples such as the van der Pol, Rayleigh, and Duffing oscillators [23]. We show numerically that, for strong dissipation and typical parameters, the steady state density matrix in the excitation number basis exhibits a power-law tail in the diagonal elements (which correspond to the probability distribution), and also in the off-diagonal elements (which correspond to the quantum coherences) with the same scaling behavior.
$M$-boson model.—
@@
Figure 1: (a) A quantum particle trapped in a one-dimensional harmonic potential interacts with $M$ independent baths ($M=2$ depicted). The particle exchanges $m$ excitation quanta with the $m$-th bath ($m=1,\ldots,M$), at a dissipation rate of $\gamma_{m}$ and pumping rate of $\kappa_{m}$. (b) Simulated measurement for the number of excitation quanta $n$ on the steady state of Eq. (1) with $M=2$ and $500$ measurement samples. The dashed line indicates the median count $n=1$, and the dotted line indicates the $90^{\text{th}}$ percentile count $n\approx 41.1$. Extreme events reaching $n\approx 16000$ are observed. Parameters are $\{\gamma_{1},\kappa_{1},\gamma_{2},\kappa_{2}\}=\{6,0,1,1\}$. (c) The probability distribution $\rho_{n,n}$ for $n$ exhibits a power-law tail $\sim n^{-\nu}$ at the critical point $\kappa_{2}=\gamma_{2}$, plotted for $\gamma_{1}/\gamma_{2}=6$ and $\kappa_{1}/\gamma_{2}=0$. The inset shows the power-law exponent $\nu$ against $\gamma_{1}/4\gamma_{2}$. The data points are obtained by numerically fitting with a power law, which agrees excellently with the analytical result  (3), $\nu=\gamma_{1}/4\gamma_{2}$. Away from the critical point, the probability distribution follows a power law up to an exponential cutoff at $n\approx 2/\delta$.
@@
Consider the system to be a quantum harmonic oscillator, with a characteristic energy scale $\hbar\omega_{0}$. The system interacts with $M$ independent baths, by exchanging $m$ excitation quanta with the $m$-th bath ($m=1,\ldots,M$), at dissipation and pumping rates denoted by $\gamma_{m}$ and $\kappa_{m}$ respectively [see Fig. 1(a)]. For a bath at thermal equilibrium, the ratio $0\leq\kappa_{m}/\gamma_{m}\leq 1$ is determined by the temperature of the bath, where the lower and upper limits correspond to zero and infinite temperatures respectively. Under the Born-Markov approximation, the quantum dynamics of the system can be modelled by the Lindblad master equation [14, 15, 16]
@@
for the density matrix $\rho$ of the system, where $\mathcal{D}[\hat{c}]\rho\equiv\hat{c}\rho\hat{c}^{\dagger}-(\hat{c}^{\dagger}\hat{c}\rho+\rho\hat{c}^{\dagger}\hat{c})/2$ is the dissipator. The annihilation and creation operators $\hat{a}$ and $\hat{a}^{{\dagger}}$ for the system excitation satisfy the bosonic commutation relation $[\hat{a},\hat{a}^{{\dagger}}]=1$. The model can physically describe, for instance, photonic excitations (in optical or microwave cavities) or phononic excitations (in a trapped ion or nanomechanical resonator). In Eq. (1), the bath degrees of freedom are traced out, and the effects of the microscopic system-bath interactions are captured by the dissipator terms. We shall refer to Eq. (1) as the $M$-boson model. The case of $M=1$ reduces to the well-known example of a damped harmonic oscillator [14, 15, 16], while the $M=2$ case was previously studied in the context of quantum limit cycles and synchronization [24, 25, 26, 27, 18]. We remark that $\gamma_{2}$ dissipation term was implemented experimentally on superconducting circuits using dissipation engineering, to stabilize Schödinger cat states for bosonic error correction [28, 29]. A similar setup can be used to implement the $\kappa_{2}$ pumping term.
Since Eq. (1) is rotationally symmetric (i.e., invariant under the $\text{U}(1)$ transformation $\hat{a}\to e^{i\phi}\hat{a}$), the steady state density matrix is diagonal in the excitation number basis {|$\{\ket{n}\}$ and independent of $\omega_{0}$. In this basis, Eq. (1) simplifies to
@@
for large $n$, small $\delta\equiv\gamma_{M}/\kappa_{M}-1\geq 0$, and any $M\geq 2$, see Supplementary Materials (SM) for details [30]. The power-law tail is exact at the critical point $\delta=0$, which can be physically interpreted as fixing the $M$-th bath at infinite temperature. The power-law behavior is almost entirely controlled by the leading order processes involving $M$ and $M-1$ bosons, which is intuitive since these processes dominate the dynamics at large $n$. By tuning the rates $\gamma_{M-1}$ and $\kappa_{M-1}$ (for example, by adjusting the temperature of the $(M-1)$-th bath), the power-law exponent $\nu$ can vary between 00 and $\infty$. The regime $\nu\leq 2$ is particularly interesting, where the mean excitation number is infinite at criticality. As shown in Fig. 1(b), this can give rise to extreme events where the measured excitation numbers are orders of magnitude higher than typical values, following the power-law behavior as shown in Fig. 1(c). More generally, the $k$-th moment of the number distribution $\rho_{n,n}$ is infinite or undefined if $\nu\leq k+1$. Away from the critical point, Eq. (3) predicts an exponential cutoff at $n\sim M/\delta$. For $\delta\ll 1$, the power-law behavior holds for a large range of $n$. This suggests that the dramatic effects of the power-law tail, such as extreme events, can be observed close to the critical point.
Figure 2: Physical signatures of the power-law tail in the $M$-boson model, plotted for $M=2$ and $\kappa_{1}=0$: (a) Mean number of excitation quanta $\braket{\hat{a}^{{\dagger}}\hat{a}}$, (b, c) Normalized second-order correlation function $g^{(2)}(0)$. $\delta=1-\kappa_{2}/\gamma_{2}$ is the deviation from the critical point. The power-law exponent is $\nu=\gamma_{1}/4\gamma_{2}$ (3). In all the plots, the solid curves are obtained numerically for $\delta\in[10^{-4},10^{-2}]$. The Hilbert space truncation is set at $N_{T}=5\times 10^{5}$ to ensure numerical convergence. The black dashed lines indicate the corresponding analytical results (5) and (7). Grey shaded regions denote the regime in which the physical quantity is divergent near the critical point, with the divergence derived from the scaling theory indicated. At large $\nu$, $g^{(2)}(0)$ scales linearly with $\nu$.
The power-law behavior at criticality might appear surprising, since the $M$-boson exchanges in and out of the system occur at precisely the same rate $\gamma_{M}=\kappa_{M}$. The emergence of the power law can be understood by examining the quantum fluctuations. Taking the white-noise limit of the bath, the master equation (1) yields, in the Heisenberg picture, the quantum Itô equation [14, 31, 17, 18]
@@
The final term showcases the quantum noise explicitly, via the independent operator-valued Wiener increments $d\hat{W}_{m}$ which satisfy $d\hat{W}_{m}d\hat{W}_{m^{\prime}}^{\dagger}=\delta_{mm^{\prime}}\gamma_{m}dt$ and $d\hat{W}_{m}^{\dagger}d\hat{W}_{m^{\prime}}=\delta_{mm^{\prime}}\kappa_{m}dt$. The noise operators are necessary to preserve the bosonic commutation relation. Crucially, the noise is additive only for $m=1$, and multiplicative for all $m>1$. In the latter case, the noise operator $d\hat{W}_{m}$ is coupled to the bosonic field $\hat{a}^{{\dagger}m-1}$. This causes the fluctuations to be self-amplified and grow rapidly, which must be counterbalanced by the damping forces for the steady state to exist. At large excitation numbers, the dominant damping term in Eq. (4) is $\propto(\gamma_{M}-\kappa_{M})\hat{a}^{{\dagger}M-1}\hat{a}^{M}$, which vanishes at the critical point $\kappa_{M}/\gamma_{M}=1$. The remaining damping forces can only weakly suppress the fluctuations at large $n$, which results in the power law tail.
Physical signatures of the power-law tail.— The power-law distribution can be directly observed in experiments by measuring the system in the excitation number basis, such as by photodetection, if the bosonic operator $\hat{a}$ represents an electromagnetic mode. Here, we explore several other physical properties in which the power-law behavior manifests. We restrict the discussion to $M=2$, which captures most of the essential physics, while analogous results for $M>2$ are discussed in the SM.

### v2

# Power-law distributions in nonequilibrium open quantum systems

@@ Abstract
Power-law probability distributions are widely used to model extreme statistical events in complex systems, with applications to a vast array of natural phenomena ranging from earthquakes to stock market crashes to pandemics. We show that analogous heavy tails arise naturally in open quantum systems with nonlinear dissipation. Introducing a prototypical family of quantum dynamical models, we analytically prove the emergence of power-law tails in the steady state energy distribution, originating from an amplification of quantum noise whose microscopic fluctuations grow with energy. Moreover, our analysis suggests a general mechanism for heavy-tail statistics in the nonequilibrium steady states of open quantum systems: Nonlinear dissipation generically induces multiplicative quantum noise, enforced by the constraints of quantum mechanics, which is responsible for the heavy-tail behavior. This is supported by numerical simulations of a general class of nonlinear dynamics known as quantum Liénard systems. Remarkably, even when the corresponding classical system is stable, we find power-law tails in both steady-state populations and coherences, which occur for typical parameters without fine-tuning. This phenomenon can potentially be harnessed to develop extreme photon sources for novel applications in light-matter interaction and sensing.
1 Introduction
@@
The natural occurrence of power-law probability distributions underpins many fascinating statistical phenomena in complex systems across vast domains such as physics, biology, geology, economics and social science [1, 2, 3, 4]. Power-law distributions are often associated with the presence of extreme or ‘black swan’ events, which are statistical events that occur with low frequency but have enormous impact [5, 6]. This is sometimes stated in various contexts as the Pareto principle: a majority of the observed effects arise from a minority of the causes. Extreme events abound in nature, with well-known examples including earthquakes [7], rogue ocean waves [8], solar flares [9], stock market crashes [10] and pandemics [11].
@@
Often, extreme events are observed as emergent phenomena in large-scale systems, and statistical models are constructed to explain the empirical data a posteriori. Here, we show that such a phenomenon can emerge dynamically in arguably one of the most fundamental stochastic systems in nature — a single dissipative quantum system. While extreme events and heavy-tail statistics (such as power laws) are well-studied in classical systems, their occurrences in quantum-mechanical systems have been largely unexplored. To this end, we construct the $M$-boson model, a prototypical toy model of a quantum dynamical system with nonlinear dissipation, and prove analytically that the energy distribution in the steady state exhibits a power-law tail. If we regard the system physically as a quantum particle trapped in a potential well, this phenomenon manifests as a power-law decay in the position and momentum distribution of the quantum particle [12, 13].
@@
The physical origin of the power-law distribution in our model can be attributed to quantum noise [14]. The phenomenon stems from the intimate connection between microscopic fluctuations and dissipation in quantum mechanics, and traces back to Heisenberg’s uncertainty principle. From the perspective of open quantum systems, the microscopic interactions between the system and the external bath, which are responsible for the dissipation, unavoidably inject noise into the system [14, 15, 16]. Furthermore, the form of this quantum noise is highly dependent on the type of dissipation. For nonlinear dissipation, this noise is typically multiplicative, which means the noise is coupled to the quantum state of the system [17, 18]. This coupling induces a self-amplification of the noise, in which the scale of the fluctuations grows with the excitation of the system. The amplified fluctuations at large excitation are only weakly suppressed by the dissipation, and the net result is a power-law distribution.
Our analysis suggests a general mechanism for power laws to emerge in the nonequilibrium steady states of open quantum systems describing quantum particles subject to dissipative forces, arising from the aforementioned multiplicative quantum noise. We provide numerical evidence for this hypothesis by studying the open quantum systems obtained via quantizing a general classical nonlinear dynamical system [19]. Of particular interest is the Liénard system, a general class of dynamical systems encompassing many paradigmatic examples such as the van der Pol, Rayleigh, and Duffing oscillators [20]. Analogous to their classical counterparts, quantum Liénard systems have been established to exhibit limit cycles [21, 22, 23, 19], and display interesting nonlinear phenomena such as synchronization, relaxation oscillations, and excitable dynamics. We numerically demonstrate a new phenomenon unique to the quantum system: for typical parameters in the classically stable regime (without fine-tuning), the steady state density matrix of quantum Liénard systems exhibits power-law tails in both the populations (i.e., diagonal elements) and the coherences (i.e., off-diagonal elements) in the Fock basis, highlighting the essential role of quantum noise.
@@
Interestingly, both the $M$-boson and quantum Liénard models can exhibit a regime where the power-law tail vanishes so slowly such that the quantum state, while normalizable, contains an infinite number of excitation quanta on average. Moreover, the distribution for the excitation number is highly skewed towards low excitation. This gives rise to extreme events in which the system, when measured in the excitation number basis, produces highly excited states containing orders of magnitude more excitation quanta than the typically measured values. This regime of the power law was previously observed experimentally in a different nonlinear quantum optical setup, which works by amplifying the vacuum fluctuations [24]. In that experiment, extreme emission events containing $\sim 10^{6}$ photons in a single pulse were observed.
@@
In the context of quantum optics, where the excitations are photons, our analytical results show that the system can exhibit an extreme case of photon superbunching in which the emitted photons are significantly more correlated in time as compared to thermal light sources [25]. The intensity correlation function (at zero delay) in this case can be made arbitrarily high, due to the strong fluctuations in the electromagnetic field. This phenomenon of power-law behavior opens up potential avenues for exploring extreme regimes of light-matter interactions [26, 27, 28], and can be a useful resource for sensing applications such as ghost imaging [29, 30].
2 $M$-boson model
@@
Figure 1: Power-law distribution in the $M$-boson model. (a) A quantum particle trapped in a one-dimensional harmonic potential interacts with $M$ independent baths ($M=2$ depicted). The particle exchanges $m$ excitation quanta with the $m$-th bath ($m=1,\ldots,M$), at a dissipation rate of $\gamma_{m}$ and pumping rate of $\kappa_{m}$. (b) Simulated measurement for the number of excitation quanta $n$ on the steady state of Eq. (1) with $M=2$ and $500$ measurement samples. The dashed line indicates the median count $n=1$, and the dotted line indicates the $90^{\text{th}}$ percentile count $n\approx 41.1$. Extreme events reaching $n\approx 16000$ are observed. Parameters are $\{\gamma_{1},\kappa_{1},\gamma_{2},\kappa_{2}\}=\{6,0,1,1\}$. (c) The probability distribution $\rho_{n,n}$ for $n$ exhibits a power-law tail $\sim n^{-\nu}$ at the critical point $\kappa_{2}=\gamma_{2}$, plotted for $\gamma_{1}/\gamma_{2}=6$ and $\kappa_{1}/\gamma_{2}=0$. The inset shows the power-law exponent $\nu$ against $\gamma_{1}/4\gamma_{2}$. The data points are obtained by numerically fitting with a power law, which agrees excellently with the analytical result (3), $\nu=\gamma_{1}/4\gamma_{2}$. Away from the critical point, the probability distribution follows a power law up to an exponential cutoff at $n\sim M/\delta$.
@@
We begin by analyzing, in detail, a minimal toy model which exhibits the power-law tail. Consider the system to be a 1D quantum harmonic oscillator, with a characteristic energy scale $\hbar\omega_{0}$. The system interacts with $M$ independent baths by exchanging $m$ excitation quanta with the $m$-th bath ($m=1,\ldots,M$), at dissipation and pumping rates denoted by $\gamma_{m}$ and $\kappa_{m}$ respectively [see Fig. 1(a)]. This can be achieved by preparing the $m$-th bath at the resonant frequency $m\omega_{0}$. For a bath at thermal equilibrium, the ratio $0\leq\kappa_{m}/\gamma_{m}\leq 1$ is determined by the temperature of the bath, where the lower and upper limits correspond to zero and infinite temperatures respectively. Under the Born-Markov approximation, the quantum dynamics of the system can be modelled by the Lindblad master equation [14, 15, 16]
@@
for the density matrix $\rho$ of the system, where $\mathcal{D}[\hat{c}]\rho\equiv\hat{c}\rho\hat{c}^{\dagger}-(\hat{c}^{\dagger}\hat{c}\rho+\rho\hat{c}^{\dagger}\hat{c})/2$ is the dissipator. The annihilation and creation operators $\hat{a}$ and $\hat{a}^{{\dagger}}$ for the system excitation satisfy the bosonic commutation relation $\hat{a},\hat{a}^{{\dagger}}]=1$. The model can physically describe, for instance, photonic (in optical or microwave cavities) or phononic excitations (in a trapped ion or nanomechanical resonator). In Eq. (1), the bath degrees of freedom are traced out, and the effects of the microscopic system-bath interactions are captured by the dissipator terms. A microscopic derivation of the $M=2$ case can be found in Refs. [31, 32, 17]; the derivation for $M>2$ follows analogously. We shall refer to Eq. (1) as the $M$-boson model. The case of $M=1$ reduces to the well-known example of a damped harmonic oscillator [14, 15, 16], while the $M=2$ case, known as the quantum Stuart-Landau oscillator  [^1] This is also commonly referred to as the quantum van der Pol oscillator in the literature, e.g., Refs. [33, 34]. However, since the corresponding classical model is the Stuart-Landau model (which describes limit-cycle oscillators near the Hopf bifurcation), this model should be more appropriately called the quantum Stuart-Landau model., was previously studied in the context of quantum limit cycles and synchronization [33, 34, 35, 36, 18]. We remark that the two-boson ($\gamma_{2}$) dissipation term was implemented experimentally on superconducting circuits using dissipation engineering, to stabilize Schrödinger cat states for bosonic quantum error correction [37, 38], and also recently in trapped ions [39, 40] to demonstrate quantum synchronization. A similar setup can be used to implement the $\kappa_{2}$ pumping term.
[…]
@@
valid for large $n$, small $\delta=1-\kappa_{M}/\gamma_{M}\geq 0$, and any $M\geq 2$ (see Appendix A for details). Since detailed balance is generally not satisfied, this describes a nonequilibrium steady state. The power-law tail is exact at the critical point $\delta=0$, which can be physically interpreted as fixing the $M$-th bath at infinite temperature. The power-law behavior is almost entirely controlled by the leading order processes involving $M$ and M1M-1 bosons, which is intuitive since these processes dominate the dynamics at large $n$. By tuning the rates $\gamma_{M-1}$ and $\kappa_{M-1}$ (for example, by adjusting the temperature of the (M1)(M-1)-th bath), the power-law exponent $\nu$ can vary between 0 and $\infty$. The regime $\nu\leq 2$ is particularly interesting, where the mean excitation number is infinite at criticality. As shown in Fig. 1(b), this can give rise to extreme events where the measured excitation numbers are orders of magnitude higher than typical values, following the power-law behavior displayed in Fig. 1(c). More generally, the $k$-th moment of the number distribution $\rho_{n,n}$ is infinite or undefined if $\nu\leq k+1$. Away from the critical point, Eq. (3) predicts an exponential cutoff at $n\sim M/\delta$. For $\delta\ll 1$, the power-law behavior holds for a large range of $n$. This suggests that the dramatic effects of the power-law tail, such as extreme events, can be observed close to the critical point. Such a scenario is also more realistic than the infinite-temperature limit $\delta=0$, where physical limitations in the setup or measurement apparatus introduce a natural high-energy cutoff in $n$.
Figure 2: Physical signatures of the power-law distribution. Numerical results for the $M$-boson model with $M=2$ and $\kappa_{1}=0$. (a) Mean number of excitation quanta $\braket{\hat{a}^{{\dagger}}\hat{a}}$, (b, c) Normalized second-order correlation function $g^{(2)}(0)$. $\delta=1-\kappa_{2}/\gamma_{2}$ is the deviation from the critical point. The power-law exponent is $\nu=\gamma_{1}/4\gamma_{2}$ [Eq. (3)]. In all the plots, the solid curves are obtained numerically for $\delta\in[10^{-4},10^{-2}]$. The Hilbert space truncation is set at $N_{T}=5\times 10^{5}$ to ensure numerical convergence. The black dashed lines indicate the corresponding analytical results (5) and (7). Grey shaded regions denote the regime in which the physical quantity is divergent near the critical point, with the divergence derived from the scaling theory indicated. At large $\nu$, $g^{(2)}(0)$ scales linearly with $\nu$.
The power-law behavior at criticality might appear surprising, since the $M$-boson exchanges in and out of the system occur at precisely the same rate $\gamma_{M}=\kappa_{M}$, and classical intuition suggests that there is no net flow of excitations. The emergence of the power law can be understood by examining the quantum fluctuations. Taking the white-noise limit of the bath, the master equation (1) yields, in the Heisenberg picture, the quantum Itô equation [14, 41, 17, 18]
@@
The final term showcases the quantum noise explicitly, via the independent operator-valued Wiener increments $d\hat{W}_{m}$ which satisfy $d\hat{W}_{m}d\hat{W}_{m^{\prime}}^{\dagger}=\delta_{mm^{\prime}}\gamma_{m}dt$ and $d\hat{W}_{m}^{\dagger}d\hat{W}_{m^{\prime}}=\delta_{mm^{\prime}}\kappa_{m}dt$. The noise operators are necessary to preserve the bosonic commutation relation $\hat{a}(t),\hat{a}^{\dagger}(t)]=1$ (in the Heisenberg picture). This is in sharp contrast with classical stochastic systems, where the noise terms can be introduced quite freely by hand (at least from a mathematical standpoint). Crucially, the quantum noise is additive only for $m=1$, and multiplicative for all $m>1$. In the latter case, the noise operator $d\hat{W}_{m}$ is coupled to the bosonic field $\hat{a}^{{\dagger}m-1}$. This causes the fluctuations to be self-amplified and grow rapidly, which must be counterbalanced by the damping forces for the steady state to exist. At large excitation numbers, the dominant damping term in Eq. (4) is $\propto(\gamma_{M}-\kappa_{M})\hat{a}^{{\dagger}M-1}\hat{a}^{M}$, which vanishes at the critical point $\kappa_{M}/\gamma_{M}=1$. The remaining damping forces can only weakly suppress the fluctuations at large $n$, which results in the power-law tail.
Our analysis reveals that the heavy-tail statistics of the $M$-boson model can be attributed to the multiplicative quantum noise in Eq. (4), which arises naturally from the nonlinear dissipation terms $\mathcal{D}[\hat{a}^{m}]\rho$ for $m>1$. More generally, for an arbitrary Lindblad operator $\hat{c}$, the noise term in Eq. (4) is of the form $\sim[\hat{c}^{\dagger},\hat{a}]d\hat{W}+[\hat{a},\hat{c}]d\hat{W}^{\dagger}$ [41], which is multiplicative if $\hat{c}$ is nonlinear in $\hat{a}$ or $\hat{a}^{{\dagger}}$. This suggests a mechanism for heavy-tail statistics to emerge in the steady state in more general nonequilibrium open quantum systems:
  1. 1.
    Quantum dissipation is invariably associated with quantum noise (this is true of any open quantum system).
  2. 2.
    Crucially, the form of the quantum noise is not arbitrary. It is heavily constrained by the type of dissipation, in order to preserve commutation relations, e.g., $\hat{a}(t),\hat{a}^{\dagger}(t)]=1$, where $\hat{a}(t)$ is the bosonic annihilation operator in the Heisenberg picture.
  3. 3.
    For nonlinear dissipation, the quantum noise is typically multiplicative. The competition between multiplicative noise (which tends to cause amplification) and nonlinear dissipation (which tends to suppress large excitations) gives rise to the heavy-tail (e.g., power-law) statistics in the steady state.
To support this argument, we will numerically study a general class of dynamical systems known as quantum Liénard systems [19] in Sec. 4, and show that power-law tails appear not only in the steady state populations $\rho_{n,n}$, but also in the steady-state coherences $\rho_{n,m}$ ($m\neq n$). Before doing so, we first discuss some measurable signatures of the power law.
3 Physical signatures of the power-law tail
The power-law distribution can be directly observed in experiments by measuring the system in the excitation number basis, such as by photodetection, if the bosonic operator $\hat{a}$ represents an electromagnetic mode. This was achieved experimentally in Ref. [24], using a different physical system based on nonlinear optics. There, the origin of the power law is also attributed to amplified quantum noise. In this section, we explore several other physical properties in which the power-law behavior manifests. For simplicity, we will illustrate this using the $M$-boson model due to its analytical tractability. However, note that much of the discussion here holds more generally for any quantum state exhibiting the power-law tail $\rho_{n,n}\sim n^{-\nu}$. We restrict to $M=2$, which already captures most of the essential physics, while analogous results for $M>2$ are discussed in the Appendix.